1. We know that the probability of correctly responding to a multiple choice question while guessing is 0.25 if there are four answer choices. Suppose a test has 20 questions. We also know that correctly responding follows a binomial distribution for this test.
what is the mean and standard deviation?

Answer by stanbon(75887)

We know that the probability of correctly responding to a multiple choice question while guessing is 0.25 if there are four answer choices. Suppose a test has 20 questions. We also know that correctly responding follows a binomial distribution for this test.
what is the mean and standard deviation?
-------------------------
mean = np = 20*0.25 = 5
std = sqrt(npq) = sqrt(5*0.75) = 1.9365
======================
Cheers,
Stan H.
===========